# SAR-KnowLIP: Towards Multimodal Foundation Models for Remote Sensing

### v1

@@ 4.2 Analysis of SAR-GEOVL-1M Text
Fig. 8(a,b) present the word frequency distributions and length statistics for target and terrain descriptions, respectively. Terrain-related prompts yield more diverse and open-form responses, reflected in denser textual distributions. Fig. 8(c,d,e) showcase representative samples from SAR-GEOVL-1M, corresponding to ground targets, marine targets, and terrain scenarios. The integration of the HCoT prompting strategy leads to a notable improvement in GPT-4.1’s reasoning performance and semantic expressiveness.
To assess the textual accuracy of SAR-GEOVL-1M, we conducted an expert evaluation on a randomly sampled 2% subset of the dataset, given its large scale. Each image was annotated with eight types of semantic information, and accuracy was independently assessed for each type. As shown in Fig. 9, the overall accuracy approaches 80%. Marine target descriptions achieve higher accuracy than ground targets, benefiting from the relatively uniform sea background. Environmental descriptions also perform well, aided by geospatial cues and lower semantic complexity. While fine-grained target attributes occasionally contain errors, the generated texts are generally accurate and coherent, offering strong support for high-level scene understanding and multimodal pre-training.
Figure 9: Accuracy evaluation of eight dimensions of information in SAR-GEOVL-1M text data.
Figure 10: TSNE feature distribution of images and two types of text. The text in SAR-GEOVL is closer to the image feature distribution, while the distribution of annotation-based text is sparse.
@@
To further compare semantic expressiveness, we selected a subset of SAR-GEOVL-1M with manual target annotations and generated templated descriptions using the method in RemoteCLIP [10]. As shown in Fig. 10, t-SNE visualization reveals that SAR-GEOVL-1M texts exhibit tighter clustering aligned with image features, whereas templated texts are more dispersed, indicating semantic sparsity. Additionally, as illustrated in Fig. 11, we compared and analyzed multiple RS multimodal datasets using two metrics: the text semantic richness index (MTLD[58]) and token count. Overall, SAR-GEOVL-1M significantly outperforms existing mainstream datasets in both semantic richness and token count.
@@ 4.3 Visual Task Benchmarks
SAR-KnowLIP ViT-Large 75.96 90.35 66.84 95.14 77.46 98.02 68.14 82.54 68.97 79.13 38.41 70.62 46.29
1) Target Classification: The classifier is implemented using the PyTorch framework and evaluated on three representative SAR datasets: FUSAR-SHIP [19], FUSAR-AIR [51], and SAR-ACD [18]. FUSAR-SHIP and FUSAR-AIR are curated subsets of SAR-GEOVL-1M, focusing on ship and aircraft target recognition, respectively. FUSAR-SHIP, developed by the Key Laboratory of Electromagnetic Wave Information Science at Fudan University, includes 5,242 images spanning 15 ship categories and 98 subcategories. FUSAR-AIR comprises diverse aircraft types such as transport, refueling, and civil aircraft. SAR-ACD, a public third-party dataset with higher scene complexity and 1-meter resolution, contains 3,032 images across six aircraft types, with approximately 500 samples per class.
@@
As shown in Table IV, the proposed SAR-KnowLIP foundation model consistently outperforms all baselines in aircraft and ship classification tasks, demonstrating superior recognition capability. On the SAR-ACD dataset, it achieves a Top-3 accuracy of 99.84%, highlighting its strength in fine-grained category discrimination. While SAR-JEPA, specifically designed for SAR classification, also delivers strong performance. MoCo v3 and MAE, both pre-trained on SAR data, outperform most multimodal models, highlighting the benefits of modality-specific pretraining. Notably, ViT-Large underperforms due to limited SAR fine-tuning data, whereas SAR-KnowLIP maintains the best performance at equal model scale, reflecting its robustness and generalization ability.
@@
SAR-KnowLIP has the best performance.
From the target detection results presented in Table IV, the SAR-KnowLIP model demonstrates significantly superior performance in the three tasks of aircraft detection, ship detection, and few-shot detection, exhibiting excellent target perception and discrimination capabilities. In contrast, GeoRSClip, BAN, BLIP, and MoCo v3 ranked second in some tasks but showed imbalanced performance across other tasks, indicating deficiencies in task generalization and robustness.
@@
The experimental results are presented in Table IV. SAR-KnowLIP achieved the highest $OA$ and $mIoU$ on both segmentation datasets, significantly outperforming existing multimodal and unimodal methods. Among the multimodal models, performance varied greatly due to differences in modeling capabilities of RS domain knowledge. Models such as RemoteClip, GeoRSClip, CLIP, and SkyScript performed relatively well. The unimodal model SAR-JEPA performed well on FUSAR-MAP but showed a significant decline in performance on the more complex AIR-PolarSAR-Seg.
Fig. 1(a) shows the performance distribution of the models across three visual tasks. In general, SAR-KnowLIP demonstrates stronger generalization ability in typical visual task scenarios by jointly modeling image and text modalities. Fig. 12 visualizes the detection results for qualitative analysis. For the same model with different ViT sizes, we show better results. Fig. 12(a) demonstrates the detection capability of SAR under noise interference. GeoChat, VHM, BITA, and SAR-JEPA all missed targets, while other methods exhibited noticeable false alarms. Only our method successfully detected all targets. Fig. 12(b) highlights the detection capability for small targets at the edges. Only our method and MoCo v3 correctly detected these targets. Combining Fig. 1(a) and Fig. 12, SAR-KnowLIP exhibits superior performance in visual tasks.
@@ 4.4 Visual-Language Tasks Benchmarks
Compared to traditional vision tasks under closed-label systems, visual-language tasks are more open and can more comprehensively examine a model’s cross-modal understanding and reasoning capabilities. To systematically evaluate the multimodal modeling performance of SAR-KnowLIP, we established a visual-language task benchmark covering three representative tasks: image captioning, image-text retrieval, and visual question answering (VQA). These three tasks correspond to semantic generation, cross-modal matching, and cross-modal reasoning, respectively, and can test the model’s generalization performance and multi-dimensional expressive capabilities in remote sensing language understanding from different dimensions.
@@
Figure 15: The multimodal model’s responses to the VQA task, with the correct responses shown in color. Overall, SAR-KnowLIP has the best performance.
@@
SAR-KnowLIP ViT-Large 26.70 39.13 51.55 29.81 61.86 75.90 47.49 79.64 58.66 117.83 52.11 98.91 93.36 98.16 90.89 96.72 81.42
1) Image-Text Retrieval: This task evaluates the model’s cross-modal retrieval capability between SAR images and text in both directions: Image→Text and Text→Image, which is crucial for RS image retrieval. Standard $Recall@K$ metrics ($R@1$, $R@5$, $R@10$) are used to assess performance, with $txt\_{r1/r5/r10}$ and img_r1/r5/r10img\_r1/r5/r10 denoting the Top-K recall rates for text-to-image and image-to-text retrieval, respectively. The average of these six values serves as the overall retrieval metric. Quantitative and qualitative results are shown in Table VI and Fig. 14. As shown in Fig. 14, we compare the proposed method with a suboptimal model. SAR-KnowLIP has the highest matching degree for the correct option and can therefore retrieve the corresponding image more accurately.
2) Image Captioning: The Caption task aims to generate accurate descriptions of key targets and their spatial distribution in SAR images using natural language. We evaluate the quality of the image descriptions using four mainstream metrics[57]: BLEU-4, METEOR, CIDEr, and SPICE. These metrics assess the description quality from four dimensions: phrase matching, semantic consistency, content relevance, and structural integrity. Table VI presents the quantitative evaluation results for each pre-trained model on the Caption task. SAR-KnowLIP achieves leading performance across many indicators, particularly demonstrating significant advantages in CIDEr.
@@
Table VI summarizes the performance across all six tasks, while Fig. 15 presents qualitative examples of model predictions for representative questions. Experimental results show that SAR-KnowLIP consistently achieves strong performance, especially in target location reasoning and comprehensive landform identification, highlighting its robust reasoning ability in SAR-based visual question answering scenarios.
@@ 4.6 Ablation experiment
In the preceding experiments, we systematically validated the overall performance of SAR-KnowLIP across a variety of vision and multimodal tasks. However, the aggregate results do not directly reveal the specific contributions of each design component. To more clearly analyze the roles of data design and model mechanisms in performance improvement, this section further conducts ablation studies.
@@
SAR-GEOVL 92.83 93.86 93.68 96.17 86.34 85.52
SAR-GEOVL 97.18 95.80 94.61 98.36 87.43 88.39
SAR-GEOVL 98.70 96.43 97.89 99.45 89.07 93.58
@@
On typical VQA tasks, the experimental results for different configurations are shown in Table VIII. The experiments demonstrate that when trained with the SAR-GEOVL dataset we constructed, the performance of each task significantly outperforms SARLANG, validating the advanced nature of the proposed data construction method in terms of semantic richness and knowledge guidance. Additionally, both the filtering and reconstruction processes play a critical role in improving performance.
Figure 16: Visualization of model feature heatmaps after training on different datasets. SAR-KnowLIP can more accurately extract target features and distinguish background information.
Fig. 16 presents a visual comparison of features extracted by different models in the target counting task. From Fig. 16(a-b), it is evident that SAR-KnowLIP exhibits stronger resistance to interference. Fig. 16(c-e) further shows that the model trained with SARLANG suffers from noticeable target omissions, while SAR-KnowLIP accurately focuses on all target areas.
@@ 5 Conclusion
In this work, we present SAR-KnowLIP, the first general-purpose multimodal foundation model designed for SAR. To support its development, we construct SAR-GEOVL-1M, the largest SAR image–text dataset to date and the first to comprehensively preserve geographic attributes. Distinct from prior datasets, SAR-GEOVL-1M achieves spatial scale consistency, cross-platform adaptability, and hierarchical semantic representation. By further integrating HCoT and SCIO, our framework offers a systematic solution to enhance cross-modal alignment and knowledge integration.
We establish a unified SAR multimodal benchmark covering 11 representative downstream tasks across both vision and language domains, and compare SAR-KnowLIP against 14 leading foundation models. The results consistently demonstrate SAR-KnowLIP’s superior performance, particularly in challenging tasks such as object counting and land-cover classification, highlighting its ability to advance beyond perceptual-level recognition toward semantic-level cognition.

### v2

@@ 4.2 Analysis of FUSAR-GEOVL-1M Text
Fig. 8(a,b) present the word frequency distributions and length statistics for target and terrain descriptions, respectively. Terrain-related prompts yield more diverse and open-form responses, reflected in denser textual distributions. Fig. 8(c,d,e) showcase representative samples from FUSAR-GEOVL-1M, corresponding to ground targets, marine targets, and terrain scenarios. The integration of the HCoT prompting strategy leads to a notable improvement in GPT-4.1’s reasoning performance and semantic expressiveness.
To assess the textual accuracy of FUSAR-GEOVL-1M, we conducted an expert evaluation on a randomly sampled 2% subset of the dataset, given its large scale. Each image was annotated with eight types of semantic information, and accuracy was independently assessed for each type. As shown in Fig. 9, the overall accuracy approaches 80%. Marine target descriptions achieve higher accuracy than ground targets, benefiting from the relatively uniform sea background. Environmental descriptions also perform well, aided by geospatial cues and lower semantic complexity. While fine-grained target attributes occasionally contain errors, the generated texts are generally accurate and coherent, offering strong support for high-level scene understanding and multimodal pre-training.
Figure 9: Accuracy evaluation of eight dimensions of information in FUSAR-GEOVL-1M text data.
Figure 10: TSNE feature distribution of images and two types of text. The text in FUSAR-GEOVL is closer to the image feature distribution, while the distribution of annotation-based text is sparse.
@@
To further compare semantic expressiveness, we selected a subset of FUSAR-GEOVL-1M with manual target annotations and generated templated descriptions using the method in RemoteCLIP [13]. As shown in Fig. 10, t-SNE visualization reveals that FUSAR-GEOVL-1M texts exhibit tighter clustering aligned with image features, whereas templated texts are more dispersed, indicating semantic sparsity. Additionally, as illustrated in Fig. 11, we compared and analyzed multiple RS multimodal datasets using two metrics: the text semantic richness index (MTLD[62]) and token count. Overall, FUSAR-GEOVL-1M significantly outperforms existing mainstream datasets in both semantic richness and token count.
@@ 4.3 Visual Task Benchmarks
FUSAR-KLIP ViT-Large 75.96 90.35 66.84 95.14 77.46 98.02 68.14 82.54 68.97 79.13 38.41 70.62 46.29
1) Target Classification: The classifier is implemented using the PyTorch framework and evaluated on three representative SAR datasets: FUSAR-SHIP [22], FUSAR-AIR [54], and SAR-ACD [21]. FUSAR-SHIP and FUSAR-AIR are curated subsets of FUSAR-GEOVL-1M, focusing on ship and aircraft target recognition, respectively. FUSAR-SHIP, developed by the Key Laboratory of Electromagnetic Wave Information Science at Fudan University, includes 5,242 images spanning 15 ship categories and 98 subcategories. FUSAR-AIR comprises diverse aircraft types such as transport, refueling, and civil aircraft. SAR-ACD, a public third-party dataset with higher scene complexity and 1-meter resolution, contains 3,032 images across six aircraft types, with approximately 500 samples per class.
@@
As shown in Table IV, the proposed FUSAR-KLIP foundation model consistently outperforms all baselines in aircraft and ship classification tasks, demonstrating superior recognition capability. On the SAR-ACD dataset, it achieves a Top-3 accuracy of 99.84%, highlighting its strength in fine-grained category discrimination. While SAR-JEPA, specifically designed for SAR classification, also delivers strong performance. MoCo v3 and MAE, both pre-trained on SAR data, outperform most multimodal models, highlighting the benefits of modality-specific pretraining. Notably, ViT-Large underperforms due to limited SAR fine-tuning data, whereas FUSAR-KLIP maintains the best performance at equal model scale, reflecting its robustness and generalization ability.
@@
FUSAR-KLIP has the best performance.
From the target detection results presented in Table IV, the FUSAR-KLIP model demonstrates significantly superior performance in the three tasks of aircraft detection, ship detection, and few-shot detection, exhibiting excellent target perception and discrimination capabilities. In contrast, GeoRSClip, BAN, BLIP, and MoCo v3 ranked second in some tasks but showed imbalanced performance across other tasks, indicating deficiencies in task generalization and robustness.
@@
The experimental results are presented in Table IV. FUSAR-KLIP achieved the highest $OA$ and $mIoU$ on both segmentation datasets, significantly outperforming existing multimodal and unimodal methods. Among the multimodal models, performance varied greatly due to differences in modeling capabilities of RS domain knowledge. Models such as RemoteClip, GeoRSClip, CLIP, and SkyScript performed relatively well. The unimodal model SAR-JEPA performed well on FUSAR-MAP but showed a significant decline in performance on the more complex AIR-PolarSAR-Seg.
Fig. 1(a) shows the performance distribution of the models across three visual tasks. In general, FUSAR-KLIP demonstrates stronger generalization ability in typical visual task scenarios by jointly modeling image and text modalities. Fig. 12 visualizes the detection results for qualitative analysis. For the same model with different ViT sizes, we show better results. Fig. 12(a) demonstrates the detection capability of SAR under noise interference. GeoChat, VHM, BITA, and SAR-JEPA all missed targets, while other methods exhibited noticeable false alarms. Only our method successfully detected all targets. Fig. 12(b) highlights the detection capability for small targets at the edges. Only our method and MoCo v3 correctly detected these targets. Combining Fig. 1(a) and Fig. 12, FUSAR-KLIP exhibits superior performance in visual tasks.
@@ 4.4 Visual-Language Tasks Benchmarks
Compared to traditional vision tasks under closed-label systems, visual-language tasks are more open and can more comprehensively examine a model’s cross-modal understanding and reasoning capabilities. To systematically evaluate the multimodal modeling performance of FUSAR-KLIP, we established a visual-language task benchmark covering three representative tasks: image captioning, image-text retrieval, and visual question answering (VQA). These three tasks correspond to semantic generation, cross-modal matching, and cross-modal reasoning, respectively, and can test the model’s generalization performance and multi-dimensional expressive capabilities in remote sensing language understanding from different dimensions.
@@
Figure 15: The multimodal model’s responses to the VQA task, with the correct responses shown in color. Overall, FUSAR-KLIP has the best performance.
@@
FUSAR-KLIP ViT-Large 26.70 39.13 51.55 29.81 61.86 75.90 47.49 79.64 58.66 117.83 52.11 98.91 93.36 98.16 90.89 96.72 81.42
1) Image-Text Retrieval: This task evaluates the model’s cross-modal retrieval capability between SAR images and text in both directions: Image→Text and Text→Image, which is crucial for RS image retrieval. Standard $Recall@K$ metrics ($R@1$, $R@5$, $R@10$) are used to assess performance, with $txt\_{r1/r5/r10}$ and img_r1/r5/r10img\_r1/r5/r10 denoting the Top-K recall rates for text-to-image and image-to-text retrieval, respectively. The average of these six values serves as the overall retrieval metric. Quantitative and qualitative results are shown in Table VI and Fig. 14. As shown in Fig. 14, we compare the proposed method with a suboptimal model. FUSAR-KLIP has the highest matching degree for the correct option and can therefore retrieve the corresponding image more accurately.
2) Image Captioning: The Caption task aims to generate accurate descriptions of key targets and their spatial distribution in SAR images using natural language. We evaluate the quality of the image descriptions using four mainstream metrics[61]: BLEU-4, METEOR, CIDEr, and SPICE. These metrics assess the description quality from four dimensions: phrase matching, semantic consistency, content relevance, and structural integrity. Table VI presents the quantitative evaluation results for each pre-trained model on the Caption task. FUSAR-KLIP achieves leading performance across many indicators, particularly demonstrating significant advantages in CIDEr.
@@
Table VI summarizes the performance across all six tasks, while Fig. 15 presents qualitative examples of model predictions for representative questions. Experimental results show that FUSAR-KLIP consistently achieves strong performance, especially in target location reasoning and comprehensive landform identification, highlighting its robust reasoning ability in SAR-based visual question answering scenarios.
@@ 4.6 Ablation experiment
In the preceding experiments, we systematically validated the overall performance of FUSAR-KLIP across a variety of vision and multimodal tasks. However, the aggregate results do not directly reveal the specific contributions of each design component. To more clearly analyze the roles of data design and model mechanisms in performance improvement, this section further conducts ablation studies.
@@
FUSAR-GEOVL 92.83 93.86 93.68 96.17 86.34 85.52
FUSAR-GEOVL 97.18 95.80 94.61 98.36 87.43 88.39
FUSAR-GEOVL 98.70 96.43 97.89 99.45 89.07 93.58
@@
On typical VQA tasks, the experimental results for different configurations are shown in Table VIII. The experiments demonstrate that when trained with the FUSAR-GEOVL dataset we constructed, the performance of each task significantly outperforms SARLANG, validating the advanced nature of the proposed data construction method in terms of semantic richness and knowledge guidance. Additionally, both the filtering and reconstruction processes play a critical role in improving performance.
Figure 16: Visualization of model feature heatmaps after training on different datasets. FUSAR-KLIP can more accurately extract target features and distinguish background information.
Fig. 16 presents a visual comparison of features extracted by different models in the target counting task. From Fig. 16(a-b), it is evident that FUSAR-KLIP exhibits stronger resistance to interference. Fig. 16(c-e) further shows that the model trained with SARLANG suffers from noticeable target omissions, while FUSAR-KLIP accurately focuses on all target areas.
@@ 5 Conclusion
In this work, we present FUSAR-KLIP, the first general-purpose multimodal foundation model designed for SAR. To support its development, we construct FUSAR-GEOVL-1M, the largest SAR image–text dataset to date and the first to comprehensively preserve geographic attributes. Distinct from prior datasets, FUSAR-GEOVL-1M achieves spatial scale consistency, cross-platform adaptability, and hierarchical semantic representation. By further integrating HCoT and SCIO, our framework offers a systematic solution to enhance cross-modal alignment and knowledge integration.
We establish a unified SAR multimodal benchmark covering 11 representative downstream tasks across both vision and language domains, and compare FUSAR-KLIP against 14 leading foundation models. The results consistently demonstrate FUSAR-KLIP’s superior performance, particularly in challenging tasks such as object counting and land-cover classification, highlighting its ability to advance beyond perceptual-level recognition toward semantic-level cognition.